Alternative Capital Investments

The investment committee of Sentry Insurance Co. is evaluating two projects, office expansion and upgrade to computer servers. The projects have different useful lives, but each requires an investment of $1,343,000. The estimated net cash flows from each project are as follows:

Net Cash Flow
Year     

Office Expansion

    

Server

1 $338,000 $446,000
2 338,000 446,000
3 338,000 446,000
4 338,000 446,000
5 338,000
6 338,000

The committee has selected a rate of 10% for purposes of net present value analysis. It also estimates that the residual value at the end of each project's useful life is $0, but at the end of the fourth year, the office expansion's residual value would be $423,000.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Required:

If required, use the minus sign to indicate a negative net present value.

1. For each project, compute the net present value. Use the present value of an annuity of $1 table above. Ignore the unequal lives of the projects. If required, round to the nearest dollar.

Office Expansion Server Upgrade
Present value of annual net cash flows $ $
Less amount to be invested $ $
Net present value $ $

2. For each project, compute the net present value, assuming that the office expansion is adjusted to a four-year life for purposes of analysis. Use the present value of $1 table above.

Office Expansion Server Upgrade
Present value of net cash flow total $ $
Less amount to be invested $ $
Net present value $ $

3. The net present value of the two projects over equal lives indicates that the OFFICE EXPANSION / SERVER UPGRADE  has a higher net present value and would be a superior investment.
Answer #1
1
Office Expansion Server Upgrade
Present value of annual net cash flows 1471990 1413820
Less amount to be invested 1343000 1343000
Net present value 128990 70820
2
Office Expansion Server Upgrade
Present value of annual net cash flows 1360031 1413374
Less amount to be invested 1343000 1343000
Net present value 17031 70374
3
The net present value of the two projects over equal lives indicates that the SERVER UPGRADE has a higher net present value and would be a superior investment.
Workings:
1
Office Expansion Server Upgrade
Present value of annual net cash flows =338000*4.355 =446000*3.170
2
Office Expansion Year 4 cash flow 761000 =338000+423000
Office Expansion Server Upgrade
Cash flows PV factor Present value Cash flows PV factor Present value
1 338000 0.909 307242 446000 0.909 405414
2 338000 0.826 279188 446000 0.826 368396
3 338000 0.751 253838 446000 0.751 334946
4 761000 0.683 519763 446000 0.683 304618
Total 1360031 Total 1413374